Find the det(-2A) and compare it to det(A) for A={1 1}{3 -1}


1
Expert's answer
2021-06-15T09:16:15-0400

Solution:-

solving det(A)

we get

"\left|\begin{array}{cc}1 & 1\\3 & -1\end{array}\right| = \left(1\right)\cdot \left(-1\right) - \left(1\right)\cdot \left(3\right) = -4" ................(1)

we know the property of determinant

"det(xA)=" "x^ndet(A)" where n is the order .

here

"det(-2A)=" "(-2)^2det(A)"

"=4(-4)" ...............(from 1)

"=-16"